# Bayesian Inference of Gravity through Realistic 3D Modeling of Wide Binary Orbits: General Algorithm and a Pilot Study with HARPS Radial Velocities

### v4

@@ 3.1 Distribution of the Bayesian Inferred Parameters
For the Newtonian model, the values obtained by Saglia:2025 are also exhibited. The Saglia:2025 values are qualitatively in agreement with the Bayesian results. However, the Saglia:2025 “best-fit” values have larger scatters (with respect to the prior expectations) than the combined Bayesian distribution because the latter takes into account individual distributions themselves. The Bayesian distributions of $e$ and $\phi_{0}$ overall follow their prior distributions well in both gravity models. However, it does not appear that $\Delta\phi$ follows well the prior expectation that is obtained by convolving Equation (13) with the $e$-values shown in the leftmost column. This is due to some surplus near one of the tails that was already noted by Saglia:2025. In the generalized gravity model, the Bayesian distribution of $\Delta\phi$ follows the prior distribution well. To estimate the statistical significance of the difference in the distribution of $\Delta\phi$ between the two gravity models, a likelihood-based comparison is considered in the following. A goodness-of-fit test of each distribution may not be so meaningful here because the Bayesian results were produced with the priors. Such a test will be considered below for the case without priors.
@@
Usually, BIC is calculated using the maximum likelihood estimates (MLE) of the parameters. Following this convention, the sorted quantities $\Delta\phi_{k}$ may be obtained from the individual MLE values as described in Appendix A. However, the MLE values are not likely to be optimal in the present case because the observational constraints are not sufficient (in particular, the radial separation between the two stars cannot be used for the MLE estimate: see Appendix A). Thus, the sorted quantities $\Delta\phi_{k}$ are obtained from the combined distribution of all Bayesian PDFs of $\Delta\phi$, so that all posterior probabilities are taken into account. Specifically, they are derived as follows. From $2\times 10^{6}$ values of $\Delta\phi$ for each of the $n=32$ wide binaries we have a total of $64\times 10^{6}$ values for the sample. These values are sorted in the ascending order to define the combined PDF. When the numerical PDF is split into $n$ bins so that each bin has $2\times 10^{6}$ values, each bin occupies one occurrence in the ascending order and thus $\Delta\phi_{k}$ is defined by the maximum value in the $k$-th bin, [^2] In the last bin, the maximum value can get arbitrarily close to the upper limit of the entire range of $\Delta\phi$ with $1-F(\Delta\phi_{n})\rightarrow 0$. Because there can arise a computer numerical error when $\sigma_{k}^{2}=nF(\Delta\phi_{k})(1-F(\Delta\phi_{k}))$ is too close to zero, I use the 99.9th percentile value in each bin rather than the maximum value. This prescription has no impact on the likelihood but can bias the value of $A_{n}^{2}$ (Equation (16)) just by $\approx-0.1$. as one can verify with mock data generated with a PDF (the rightmost bottom panel provides such an example). BIC values are calculated for both the Newtonian and the generalized gravity results, and their difference $\Delta\rm{BIC}_{\Delta\phi}={\rm BIC}_{\Delta\phi,{\rm Newton}}-{\rm BIC}_{\Delta\phi,{\rm general}}$ is given in the top panel. The value of $\Delta{\rm BIC}_{\Delta\phi}=65.5$ indicates a very strong preference for the generalized gravity model based on the conventional criterion $\Delta\rm{BIC}>10$ (Jeffreys:1939; Kass:1995). If MLE values of $\Delta\phi_{k}$ are used, a somewhat larger value of $\Delta{\rm BIC}_{\Delta\phi}=83.9$ is obtained. If the Saglia:2025 best-fit values are used, we have a somewhat smaller value of $\Delta{\rm BIC}_{\Delta\phi}=39.1$ which is still very significant. Throughout I will use only the above standard estimates of $\Delta\phi_{k}$ based on the PDFs unless otherwise indicated.
@@
To investigate the origin of the above large value of $\Delta\rm{BIC}_{\Delta\phi}$ in the context of testing gravity, two subsamples with $g_{\rm N}<10^{-9}$ m s-2 (9 binaries) or $g_{\rm N}>10^{-9}$ m s-2 (23 binaries) are considered. The subsamples have $\Delta\rm{BIC}_{\Delta\phi}=25.5$ ($10^{-9}$) and $14.5$ ($10^{-9}$). Although the low-acceleration subsample is much smaller than the higher-acceleration subsample, its $\Delta\rm{BIC}_{\Delta\phi}$ value is significantly larger. In detail, the low-acceleration subsample alone strongly disfavors Newton with >10>10. However, the higher-acceleration subsample also has a nominal value somewhat larger than 10. This may be interpreted as an indication that there may be some moderate problems with standard gravity near $10^{-9}$ (e.g., $10^{-8.5}<g_{\rm N}<10^{-9}$: see Figure 1 of Chae:2024a), and/or the flexibility of the generalized model with free $\Gamma$ accounts for the uncertainty of the individual mass compared with a fixed gravity model. I also note that the values of $\Delta\rm{BIC}_{\Delta\phi}$ for the two subsamples do not add up to the value for the total sample, meaning that an increase in sample size enhances statistical significance more steeply than linearly. This is in line with the generic Kolmogorov-Smirnov (K-S) statistics for CDF data.
@@
To further test Newtonian gravity in an absolute sense, the Anderson-Darling (A-D; AndersonDarling:1954) test is considered here since it is more sensitive to the data at the tails than the K-S test. The A-D statistic is
@@
where $F(\Delta\phi)$ is the CDF given by Equation (14). The Newtonian model has $A_{n}^{2}=5.4$ with a $p$-value (Marsaglia:2004) of $1.9\times 10^{-3}$ (or $\approx 3.1\sigma$ significance) which means a significant discrepancy with the $e$-convolved distribution $f_{e\text{-conv}}(\Delta\phi)$. The MOND-type toy model has $A_{n}^{2}=4.9$ with a $p$-value of $3.3\times 10^{-3}$ ($\approx 2.9\sigma$ significance). Based on a contemporary criterion of $p<5\times 10^{-3}$ (Wasserstein:2016; Wasserstein:2019) or $3\sigma$ for significance, the Newtonian model is significantly discrepant with the prior expectation while the significance is relatively weak for the MOND-type toy model although it appears not to be a good model either. Thus, in line with the model comparison with $\Delta{\rm BIC}_{\Delta\phi}$, Newtonian gravity has a significant discrepancy with the data in an absolute sense, although the discrepancy is not a $5\sigma$ level given the small sample size. This pilot analysis of the Bayesian inferred parameters demonstrates that a much larger sample can more decisively test gravity in the future.
@@
All fitted parameters and some derived parameters for Newtonian and generalized gravities are listed in Tables 1 and 2. In not only generalized gravity but also Newtonian gravity, the fitted parameters in general have broad ranges due to the measurement uncertainties. Different estimates of eccentricity and gravity parameters are compared in Figure 7. The Saglia:2025 values and the Bayesian results of $e$ and $\log_{10}g_{\rm N}$ are compared in the left panels of Figure 7. Here the Bayesian results refer to those obtained with priors on $e$, $i$, and $\Delta\phi$. There are reasonable agreements between the two calculations, except for one system #22, which is marked by a red circle in the left panels: Stars 36OphB and GJ663A. Without system #22, the Pearson correlation coefficient and the $p$-value between the Saglia:2025 values and the Bayesian medians are $0.43$ and $1.5\times 10^{-2}$ for $e$, and $0.94$ and $2.7\times 10^{-15}$ for $\log_{10}g_{\rm N}$. The binary system #22 has a small projected separation of s=0.030s=0.030kau (0.000140.00014pc), perhaps similar to the Sun-Halley system. Saglia:2025 obtain $r\approx a\approx 1.2$ kau that is about 40 times $s$, an unusually large value. Consequently, the predicted Newtonian gravity has a low value of $\log_{10}g_{\rm N}(\text{m\,s}^{-2})=-8.24$. On the other hand, the Bayesian output with priors predicts $r=0.07_{-0.03}^{+0.20}$kau, $a=0.03_{-0.02}^{+0.11}$, and $\log_{10}g_{\rm N}(\text{m\,s}^{-2})=-5.7_{-1.2}^{+0.6}$. The dramatic difference stems from the fact that Saglia:2025 predict a low eccentricity of $e=0.25$ while this work predicts $e=0.96_{-0.14}^{+0.02}$ similar to that of Halley’s comet. However, the Bayesian output with no/flat priors on $e$, $i$, and $\Delta\phi$ predicts $e=0.66_{-0.36}^{+0.31}$, $r=0.54_{-0.50}^{+0.85}$kau, $a=0.34_{-0.32}^{+1.12}$, and $\log_{10}g_{\rm N}(\text{m\,s}^{-2})=-7.5_{-0.8}^{+2.2}$. Because these ranges of the parameters are so broad, they are consistent with both the Bayesian results with priors and the Saglia:2025 values. However, the median values of $a$ and $r$ are perhaps too large compared with s=0.030s=0.030kau.
Table 1: Summary of Fitted and Derived Parameters in Newtonian Gravity ($\Gamma=0$) for the sample of Saglia:2025
@@
a Gaia DR3 identifications can be found in the data file given at Zenodo(Chae:Zenodo2025). b In these systems, the ordering of stars A and B is opposite to that of stars 1 and 2 from Saglia:2025 because this work follows the rule that star A is brighter (although this is dubious when the two stars are equally bright). c Corrected from the Saglia:2025 identifier “G56.3B”. d Corrected from the Saglia:2025 identifier “CPD-5211628B”. e There exists a degenerate parameter space localized around the periastron $\Delta\phi=360^{\circ}$ that is nearly indistinguishable from the solution given here. The prior probability distribution given by Equation (13) clearly prefers the latter. Also, the solution near the periastron predicts $a=2.88_{-0.99}^{+2.35}$ kau for the semimajor axis which is much larger than both the value given here and the sky-projected separation of $0.196$ kau.
@@
The right panels of Figure 7 exhibit individual values of $e$ and $\Gamma$ in generalized gravity. Here, the results with the general 3D model described in this work and the simple 3D model of Chae:2025 are compared. Two estimates of $e$ generally agree. However, there are some extreme values of $e\approx 1$ only from the simple 3D model for some systems, although the general 3D model also predicts high values >0.6>0.6 for the same systems. This shows that the reduction of one of the two orientation angles in the simple 3D model can result in somewhat biased estimates of $e$ in some individual cases (despite overall agreement in a statistical sample). However, two estimates of $\Gamma$ generally agree well without any exceptional cases. This may not be surprising, because the simple 3D model of Chae:2025 was designed to mainly measure $\Gamma$ and tested with mock data, while the inclination and angle variables are merely nuisance parameters as stated in that paper.
@@
Figure 7: The left panels show eccentricity ($e$) and Newtonian gravitational acceleration ($g_{\rm N}$) derived in Newtonian gravity. The Bayesian results with priors are compared with the Saglia:2025 numerical solutions. The one marked with red circle (binary #22: Stars 36OphB and GJ663A) exhibits a dramatic discrepancy between the two. The discrepancy is significantly reduced for the Bayesian result with no/flat priors as indicated by the blue triangle. However, the latter model overall has much less likely parameter values (see the text). The right panels show $e$ and $\Gamma$ (Equation (1)) in generalized gravity. The results from the general Bayesian methodology are compared with those from the simple Bayesian methodology of Chae:2025.
3.2 Bayesian Inference of the Gravitational Anomaly Parameter $\Gamma$
@@
While Section 3.1 is concerned with statistical comparisons and tests of gravity models based on the Bayesian PDFs of the parameters, this subsection is concerned with direct measurements of $\Gamma$ through the consolidation of the individual PDFs. For this, the logical priors (arising from the randomness of the sample) must be imposed on inclination $i$ and phase $\Delta\phi$ because of the degeneracy between gravity and these parameters. Eccentricity $e$ can also have some degeneracy with gravity in general. However, it turns out that $e$ is reasonably constrained regardless of the prior when the 3D velocity $\mathbf{v}$ is very precise as in the present sample. Also, any prior on $e$ has to be empirical rather than a logical requirement. I consider two cases, i.e., the generic thermal prior ($f_{\rm pr}(e)=2e$) and the no/flat prior.
@@
Figure 8: The individual Bayesian posterior PDFs of $\Gamma$ (Equation (1)) and their consolidated distribution are exhibited for two subsamples of binaries with $g_{\rm N}<10^{-9}$ or $g_{\rm N}>10^{-9}$m s-2. In the former, cases including or excluding Binary #24 (Stars HD189739 and HD189760) (that is gravitationally unbound by the Newtonian criterion) are considered. The result based on the simple Bayesian algorithm of Chae:2025 is also shown for the case including #24. The upper row shows the results with the thermal prior on eccentricity while the lower row is for the flat prior.
@@
Figure 8 shows the consolidated PDFs of $\Gamma$ for two subsamples with $g_{\rm N}<10^{-9}$ or $g_{\rm N}>10^{-9}$m s-2, where $g_{\rm N}$ is taken from the Bayesian results with generalized gravity (this is why the number in the low-acceleration regime is now 8, one less than the number with Newtonian gravity). Although the range $10^{-9}\lesssim g_{\rm N}\lesssim 10^{-8}$m s-2 belongs to a smooth transition regime in MOND gravity, the range $g_{\rm N}>10^{-9}$m s-2 is not split into finer ranges and referred to as the high-acceleration range because of the small sample size. The subsample of 24 binaries in the high-acceleration regime accurately reproduces Newtonian gravity with $\Gamma=-0.002_{-0.018}^{+0.012}$ (thermal prior) or $0.004_{-0.013}^{+0.013}$ (flat prior). This agrees with the result of Chae:2025 for a much larger sample of 125 binaries (in a high-acceleration regime) that have, however, larger uncertainties of RVs. These results extend the validity of standard gravity to gravitational acceleration as low as $10^{-8}$ or $10^{-9}$m s-2.
@@
In contrast, the consolidated probability distribution of $\Gamma$ in the low-acceleration regime exhibits a discrepancy with Newton: $\Gamma=0.134_{-0.036}^{+0.056}$ (thermal prior) or $0.115_{-0.028}^{+0.048}$ (flat prior). These results represent $3.7\sigma$ inconsistency with Newton $\Gamma=0$, where the quoted errors come from one half of the nominal $2\sigma$ width (rather than the nominal $1\sigma$ width) because the $2\sigma$ width better describes the overall behavior of the consolidated PDF that is asymmetrical and somewhat different from the Gaussian shape. This high statistical significance is in large part due to one system, binary #24 (Stars HD189739 and HD189760). Saglia:2025 suggest that Binary #24 is gravitationally unbound. The Bayesian result with Newtonian gravity returns $v_{\rm obs}/v_{\rm escN}=1.116\pm 0.030$ confirming that the observed relative velocity between the pair exceeds the predicted Newtonian escape velocity. Moreover, Newtonian gravity predicts a much larger value of $a=51.2_{-21.3}^{+52.1}$ kau than the observed separation of $s=5.78$ kau, an increased mass with $\log_{10}f_{M}=0.035\pm 0.019$ (i.e., $\approx 8$% increase), a phase close to the periastron ($\Delta\phi=333_{-17}^{+15}$ degrees), and a significantly shifted posterior velocity of $v_{\rm mod}=0.82\pm 0.02$ km s-1 (Equation (3)), which means $v_{\rm mod}/v_{\rm obs}=0.87\pm 0.02$ and $v_{\rm mod}/v_{\rm escN}=0.96\pm 0.03$. Thus, this system is clearly in tension with Newtonian gravity.

### v5

@@ III.1 Distribution of the Bayesian Inferred Parameters
For the Newtonian model, the values obtained by Saglia et al. (2025) are also exhibited. The Saglia et al. (2025) values are qualitatively in agreement with the Bayesian results. However, the Saglia et al. (2025) “best-fit” values have larger scatters (with respect to the prior expectations) than the combined Bayesian distribution because the latter takes into account individual distributions themselves. The Bayesian distributions of $e$ and $\phi_{0}$ overall follow their prior distributions well in both gravity models. However, it does not appear that $\Delta\phi$ follows well the prior expectation that is obtained by convolving Equation (13) with the $e$-values shown in the leftmost column. This is due to some surplus near one of the tails that was already noted by Saglia et al. (2025). In the generalized gravity model, the Bayesian distribution of $\Delta\phi$ follows the prior distribution well. To estimate the statistical significance of the difference in the distribution of $\Delta\phi$ between the two gravity models, a likelihood-based comparison is considered in the following. A goodness-of-fit test of each distribution may not be so meaningful here because the Bayesian results were produced with the priors. Such a test will be considered below for the case without priors.
@@
Usually, BIC is calculated using the maximum likelihood estimates (MLE) of the parameters. Following this convention, the sorted quantities $\Delta\phi_{k}$ may be obtained from the individual MLE values as described in Appendix A. However, the MLE values are not likely to be optimal in the present case because the observational constraints are not sufficient (in particular, the radial separation between the two stars cannot be used for the MLE estimate: see Appendix A). Thus, the sorted quantities $\Delta\phi_{k}$ are obtained from the combined distribution of all Bayesian PDFs of $\Delta\phi$, so that all posterior probabilities are taken into account. Specifically, they are derived as follows. From $2\times 10^{6}$ values of $\Delta\phi$ for each of the $n=32$ wide binaries we have a total of $64\times 10^{6}$ values for the sample. These values are sorted in the ascending order to define the combined PDF. When the numerical PDF is split into $n$ bins so that each bin has $2\times 10^{6}$ values, each bin occupies one occurrence in the ascending order and thus $\Delta\phi_{k}$ is defined by the maximum value in the $k$-th bin, [^3] In the last bin, the maximum value can get arbitrarily close to the upper limit of the entire range of $\Delta\phi$ with $1-F(\Delta\phi_{n})\rightarrow 0$. Because there can arise a computer numerical error when $\sigma_{k}^{2}=nF(\Delta\phi_{k})(1-F(\Delta\phi_{k}))$ is too close to zero, I use the 99.9th percentile value in each bin rather than the maximum value. This prescription has no impact on the likelihood but can bias the value of $A_{n}^{2}$ (Equation (16)) just by $\approx-0.1$. as one can verify with mock data generated with a PDF (the rightmost bottom panel provides such an example). BIC values are calculated for both the Newtonian and the generalized gravity results, and their difference $\Delta\rm{BIC}_{\Delta\phi}={\rm BIC}_{\Delta\phi,{\rm Newton}}-{\rm BIC}_{\Delta\phi,{\rm general}}$ is given in the top panel. The value of $\Delta{\rm BIC}_{\Delta\phi}=65.5$ indicates a very strong preference for the generalized gravity model based on the conventional criterion $\Delta\rm{BIC}>10$ (Jeffreys, 1939; Kass and Raftery, 1995). If MLE values of $\Delta\phi_{k}$ are used, a somewhat larger value of $\Delta{\rm BIC}_{\Delta\phi}=83.9$ is obtained. If the Saglia et al. (2025) best-fit values are used, we have a somewhat smaller value of $\Delta{\rm BIC}_{\Delta\phi}=39.1$ which is still very significant. Throughout I will use only the above standard estimates of $\Delta\phi_{k}$ based on the PDFs unless otherwise indicated.
@@
To investigate the origin of the above large value of $\Delta\rm{BIC}_{\Delta\phi}$ in the context of testing gravity, two subsamples with $g_{\rm N}<10^{-9}$ m s-2 (9 binaries) or $g_{\rm N}>10^{-9}$ m s-2 (23 binaries) are considered. The subsamples have $\Delta\rm{BIC}_{\Delta\phi}=25.5$ ($10^{-9}$) and $14.5$ ($10^{-9}$). Although the low-acceleration subsample is much smaller than the higher-acceleration subsample, its $\Delta\rm{BIC}_{\Delta\phi}$ value is significantly larger. In detail, the low-acceleration subsample alone strongly disfavors Newton with >10>10. However, the higher-acceleration subsample also has a nominal value somewhat larger than 10. This may be interpreted as an indication that there may be some moderate problems with standard gravity near $10^{-9}$ (e.g., $10^{-8.5}<g_{\rm N}<10^{-9}$: see Figure 1 of Chae 2024b), and/or the flexibility of the generalized model with free $\Gamma$ accounts for the uncertainty of the individual mass compared with a fixed gravity model. I also note that the values of $\Delta\rm{BIC}_{\Delta\phi}$ for the two subsamples do not add up to the value for the total sample, meaning that an increase in sample size enhances statistical significance more steeply than linearly. This is in line with the generic Kolmogorov-Smirnov (K-S) statistics for CDF data.
@@
To further test Newtonian gravity in an absolute sense, the Anderson-Darling (A-D; Anderson and Darling, 1954) test is considered here since it is more sensitive to the data at the tails than the K-S test. The A-D statistic is
@@
where $F(\Delta\phi)$ is the CDF given by Equation (14). The Newtonian model has $A_{n}^{2}=5.4$ with a $p$-value (Marsaglia and Marsaglia, 2004) of $1.9\times 10^{-3}$ (or $\approx 3.1\sigma$ significance) which means a significant discrepancy with the $e$-convolved distribution $f_{e\text{-conv}}(\Delta\phi)$. The MOND-type toy model has $A_{n}^{2}=4.9$ with a $p$-value of $3.3\times 10^{-3}$ ($\approx 2.9\sigma$ significance). Based on a contemporary criterion of $p<5\times 10^{-3}$ (Wasserstein and Lazar, 2016; Wasserstein et al., 2019) or $3\sigma$ for significance, the Newtonian model is significantly discrepant with the prior expectation while the significance is relatively weak for the MOND-type toy model although it appears not to be a good model either. Thus, in line with the model comparison with $\Delta{\rm BIC}_{\Delta\phi}$, Newtonian gravity has a significant discrepancy with the data in an absolute sense, although the discrepancy is not a $5\sigma$ level given the small sample size. This pilot analysis of the Bayesian inferred parameters demonstrates that a much larger sample can more decisively test gravity in the future.
@@
All fitted parameters and some derived parameters for Newtonian and generalized gravities are listed in Tables 1 and 2. In not only generalized gravity but also Newtonian gravity, the fitted parameters in general have broad ranges due to the measurement uncertainties. Different estimates of eccentricity and gravity parameters are compared in Figure 7. The Saglia et al. (2025) values and the Bayesian results of $e$ and $\log_{10}g_{\rm N}$ are compared in the left panels of Figure 7. Here the Bayesian results refer to those obtained with priors on $e$, $i$, and $\Delta\phi$. There are reasonable agreements between the two calculations, except for one system #22, which is marked by a red circle in the left panels: Stars 36OphB and GJ663A. Without system #22, the Pearson correlation coefficient and the $p$-value between the Saglia et al. (2025) values and the Bayesian medians are $0.43$ and $1.5\times 10^{-2}$ for $e$, and $0.94$ and $2.7\times 10^{-15}$ for $\log_{10}g_{\rm N}$. The binary system #22 has a small projected separation of s=0.030s=0.030kau (0.000140.00014pc), perhaps similar to the Sun-Halley system. Saglia et al. (2025) obtain $r\approx a\approx 1.2$ kau that is about 40 times $s$, an unusually large value. Consequently, the predicted Newtonian gravity has a low value of $\log_{10}g_{\rm N}(\text{m\,s}^{-2})=-8.24$. On the other hand, the Bayesian output with priors predicts $r=0.07_{-0.03}^{+0.20}$kau, $a=0.03_{-0.02}^{+0.11}$, and $\log_{10}g_{\rm N}(\text{m\,s}^{-2})=-5.7_{-1.2}^{+0.6}$. The dramatic difference stems from the fact that Saglia et al. (2025) predict a low eccentricity of $e=0.25$ while this work predicts $e=0.96_{-0.14}^{+0.02}$ similar to that of Halley’s comet. However, the Bayesian output with no/flat priors on $e$, $i$, and $\Delta\phi$ predicts $e=0.66_{-0.36}^{+0.31}$, $r=0.54_{-0.50}^{+0.85}$kau, $a=0.34_{-0.32}^{+1.12}$, and $\log_{10}g_{\rm N}(\text{m\,s}^{-2})=-7.5_{-0.8}^{+2.2}$. Because these ranges of the parameters are so broad, they are consistent with both the Bayesian results with priors and the Saglia et al. (2025) values. However, the median values of $a$ and $r$ are perhaps too large compared with s=0.030s=0.030kau.
Table 1: Summary of Fitted and Derived Parameters in Newtonian Gravity ($\Gamma=0$) for the sample of Saglia et al. (2025)
@@
a Gaia DR3 identifications can be found in the data file given at Zenodo(Chae, 2025a). b In these systems, the ordering of stars A and B is opposite to that of stars 1 and 2 from Saglia et al. (2025) because this work follows the rule that star A is brighter (although this is dubious when the two stars are equally bright). c Corrected from the Saglia et al. (2025) identifier “G56.3B”. d Corrected from the Saglia et al. (2025) identifier “CPD-5211628B”. e There exists a degenerate parameter space localized around the periastron $\Delta\phi=360^{\circ}$ that is nearly indistinguishable from the solution given here. The prior probability distribution given by Equation (13) clearly prefers the latter. Also, the solution near the periastron predicts $a=2.88_{-0.99}^{+2.35}$ kau for the semimajor axis which is much larger than both the value given here and the sky-projected separation of $0.196$ kau. [Note: The argument of ascending node is not given here and in all the other following tables because it is related to the parameter $\theta$ that can be obtained by Equation (5) from the other parameters given here.]
@@
The right panels of Figure 7 exhibit individual values of $e$ and $\Gamma$ in generalized gravity. Here, the results with the general 3D model described in this work and the simple 3D model of Chae (2025b) are compared. Two estimates of $e$ generally agree. However, there are some extreme values of $e\approx 1$ only from the simple 3D model for some systems, although the general 3D model also predicts high values >0.6>0.6 for the same systems. This shows that the reduction of one of the two orientation angles in the simple 3D model can result in somewhat biased estimates of $e$ in some individual cases (despite overall agreement in a statistical sample). However, two estimates of $\Gamma$ generally agree well without any exceptional cases. This may not be surprising, because the simple 3D model of Chae (2025b) was designed to mainly measure $\Gamma$ and tested with mock data, while the inclination and angle variables are merely nuisance parameters as stated in that paper.
@@
Figure 7: The left panels show eccentricity ($e$) and Newtonian gravitational acceleration ($g_{\rm N}$) derived in Newtonian gravity. The Bayesian results with priors are compared with the Saglia et al. (2025) numerical solutions. The one marked with red circle (binary #22: Stars 36OphB and GJ663A) exhibits a dramatic discrepancy between the two. The discrepancy is significantly reduced for the Bayesian result with no/flat priors as indicated by the blue triangle. However, the latter model overall has much less likely parameter values (see the text). The right panels show $e$ and $\Gamma$ (Equation (1)) in generalized gravity. The results from the general Bayesian methodology are compared with those from the simple Bayesian methodology of Chae (2025b).
III.2 Bayesian Inference of the Gravitational Anomaly Parameter $\Gamma$
@@
While Section III.1 is concerned with statistical comparisons and tests of gravity models based on the Bayesian PDFs of the parameters, this subsection is concerned with direct measurements of $\Gamma$ through the consolidation of the individual PDFs. For this, the logical priors (arising from the randomness of the sample) must be imposed on inclination $i$ and phase $\Delta\phi$ because of the degeneracy between gravity and these parameters. Eccentricity $e$ can also have some degeneracy with gravity in general. However, it turns out that $e$ is reasonably constrained regardless of the prior when the 3D velocity $\mathbf{v}$ is very precise as in the present sample. Also, any prior on $e$ has to be empirical rather than a logical requirement. I consider two cases, i.e., the generic thermal prior ($f_{\rm pr}(e)=2e$) and the no/flat prior.
@@
Figure 8: The individual Bayesian posterior PDFs of $\Gamma$ (Equation (1)) and their consolidated distribution are exhibited for two subsamples of binaries with $g_{\rm N}<10^{-9}$ or $g_{\rm N}>10^{-9}$m s-2. In the former, cases including or excluding Binary #24 (Stars HD189739 and HD189760) (that is gravitationally unbound by the Newtonian criterion) are considered. The result based on the simple Bayesian algorithm of Chae (2025b) is also shown for the case including #24. The upper row shows the results with the thermal prior on eccentricity while the lower row is for the flat prior.
@@
Figure 8 shows the consolidated PDFs of $\Gamma$ for two subsamples with $g_{\rm N}<10^{-9}$ or $g_{\rm N}>10^{-9}$m s-2, where $g_{\rm N}$ is taken from the Bayesian results with generalized gravity (this is why the number in the low-acceleration regime is now 8, one less than the number with Newtonian gravity). Although the range $10^{-9}\lesssim g_{\rm N}\lesssim 10^{-8}$m s-2 belongs to a smooth transition regime in MOND gravity, the range $g_{\rm N}>10^{-9}$m s-2 is not split into finer ranges and referred to as the high-acceleration range because of the small sample size. The subsample of 24 binaries in the high-acceleration regime accurately reproduces Newtonian gravity with $\Gamma=-0.002_{-0.018}^{+0.012}$ (thermal prior) or $0.004_{-0.013}^{+0.013}$ (flat prior). This agrees with the result of Chae (2025b) for a much larger sample of 125 binaries (in a high-acceleration regime) that have, however, larger uncertainties of RVs. These results extend the validity of standard gravity to gravitational acceleration as low as $10^{-8}$ or $10^{-9}$ m s-2.
In contrast, the consolidated probability distribution of $\Gamma$ in the low-acceleration regime exhibits a discrepancy with Newton: $\Gamma=0.134_{-0.036}^{+0.056}$ (thermal prior) or $0.115_{-0.028}^{+0.048}$ (flat prior). These results represent $3.7\sigma$ inconsistency with Newton $\Gamma=0$, where the quoted errors come from one half of the nominal $2\sigma$ width (rather than the nominal $1\sigma$ width) because the $2\sigma$ width better describes the overall behavior of the consolidated PDF that is asymmetrical and somewhat different from the Gaussian shape. If a stricter limit of $g_{\rm N}<10^{-9.5}$ m s-2 is considered [^4] Although this limit is still not fully satisfactory, MONDian gravity is expected to dominate at accelerations lower than this (see Figure 1 of Chae (2024b) and also Pflamm-Altenburg (2025)). For the present sample, only 3 wide binaries satisfy $g_{\rm N}<a_{0}$. for the internal acceleration to be more consistent with the MOND regime, only 6 wide binaries meet the limit. For these wide binaries, it is found that $\Gamma=0.143_{-0.041}^{+0.068}$ (thermal prior) or $0.149_{-0.044}^{+0.054}$ (flat prior), which are similar for the 8 wide binaries above.
@@
The moderately high statistical significance for $\Gamma>0$ is in large part due to one system, binary #24 (Stars HD189739 and HD189760). Saglia et al. (2025) suggest that Binary #24 is gravitationally unbound. The Bayesian result with Newtonian gravity returns $v_{\rm obs}/v_{\rm escN}=1.116\pm 0.030$ confirming that the observed relative velocity between the pair exceeds the predicted Newtonian escape velocity. Moreover, Newtonian gravity predicts a much larger value of $a=51.2_{-21.3}^{+52.1}$ kau than the observed separation of $s=5.78$ kau, an increased mass with $\log_{10}f_{M}=0.035\pm 0.019$ (i.e., $\approx 8$% increase), a phase close to the periastron ($\Delta\phi=333_{-17}^{+15}$ degrees), and a significantly shifted posterior velocity of $v_{\rm mod}=0.82\pm 0.02$ km s-1 (Equation (3)), which means $v_{\rm mod}/v_{\rm obs}=0.87\pm 0.02$ and $v_{\rm mod}/v_{\rm escN}=0.96\pm 0.03$. Thus, this system is clearly in tension with Newtonian gravity.